Is ( 3, 4, -2 ) a solution of the following system or not?


5x - y + 2z = 7

-2x + 6y + 9z = 0

-7x + 5y + 3z = -7




1
Expert's answer
2021-04-29T17:14:00-0400

Given equation are-

5x - y + 2z = 7

-2x + 6y + 9z = 0

-7x + 5y + 3z = -7


Puuting (3,4,-2) in above equation-


"\Rightarrow5(3)-4+2(-7)=11-14=-3\neq 7"


So (3,4,-2) is not the solution of above equation.